# Ultralight Polarity-Split Neuromorphic SNN for Event-Stream Super-Resolution

### v1

@@ 4.4 Ablation Study
Baseline + LearnSTPLoss 0.2696 0.1415 0.2295 97.14
Baseline + DualForward + LearnSTPLoss 0.2661 0.1404 0.2260 96.12
@@ Appendix B Supplementary Visualization Results
Data Representation. E2F methods convert the original asynchronous event stream into synchronous frame-based forms, such as event count maps, event stacks, or voxel grids [29, 30, 31]. These representations discretize the temporal domain into fixed bins, allowing the use of standard CNNs designed for images or videos. Conversely, E2E methods directly process raw event tuples $(x,y,t,p)$ and preserve the asynchronous temporal structure throughout the network, usually requiring specialized architectures such as Spiking Neural Networks (SNNs) [32].
@@ D.1 Spiking Neural Networks
Spiking Neural Networks (SNNs) are considered the third generation of neural networks, offering a biologically plausible framework for processing temporal data in a highly efficient and event-driven manner [33]. Unlike conventional artificial neural networks (ANNs) that rely on continuous-valued activations, SNNs communicate through discrete spikes, enabling sparse and asynchronous computations that closely mimic the behavior of real biological neurons. Each neuron in an SNN integrates incoming spikes over time through a membrane potential and generates an output spike only when this potential exceeds a certain threshold [34].
This temporal coding scheme allows SNNs to naturally model time-dependent dynamics, making them especially suitable for neuromorphic sensors such as event cameras, which output asynchronous streams of spatiotemporal events. Additionally, SNNs have demonstrated promising energy efficiency when deployed on neuromorphic hardware like Intel’s Loihi [35] or the SpiNNaker platform [36], due to their sparse activity and low-power operation. However, training SNNs remains challenging because the non-differentiable nature of spike generation complicates the use of standard backpropagation. To address this, surrogate gradient methods [37] and ANN-to-SNN conversion techniques [38] have been widely studied, enabling SNNs to learn complex tasks with performance approaching that of traditional deep learning models.
@@ D.2 SRM-based Spiking Neuron Encoding (Extended Explanation)
In this section, we provide a detailed description of the SRM-based spiking neuron dynamics used in our SNN architecture for event stream super-resolution [39]. Our model leverages temporal encoding and biological realism to effectively process asynchronous spike events.
@@ D.4 PSP-based Residual Upsampling Bypass
Compared to the main spiking pathway, which involves quantized spike-based transmission and nonlinear neuron thresholds, the PSP-based bypass maintains a smoother, continuous approximation of the input signal. This helps preserve fine-grained edge and texture information that may be lost during spike quantization, especially under sparse input conditions. Such a design aligns with similar principles found in residual learning [40], where shortcut connections assist in gradient propagation and retain low-level details.
@@ E.3 Downstream Application: Object Classification
As described in the main paper, in addition to image reconstruction, we also evaluated the performance of the super-resolved event data on an object classification task. We employed the Event Spike Tensor-based event classification method proposed by Gehrig et al. [41] for this evaluation. The experimental results are presented in Table 9. Since the baseline method has already achieved the highest classification accuracy, equivalent to that of original high-resolution events, on the N-MNIST and ASL-DVS datasets, our method maintains the same performance. However, on the CIFAR10-DVS dataset, our method shows a slight improvement.

### v2

@@ 4.4 Ablation Study
Baseline + w/L 0.2696 0.1415 0.2295 97.14
Baseline + DualForward + w/L 0.2661 0.1404 0.2260 96.12
@@ Appendix B Supplementary Visualization Results
Data Representation. E2F methods convert the original asynchronous event stream into synchronous frame-based forms, such as event count maps, event stacks, or voxel grids (EF1; EF3; EF6). These representations discretize the temporal domain into fixed bins, allowing the use of standard CNNs designed for images or videos. Conversely, E2E methods directly process raw event tuples $(x,y,t,p)$ and preserve the asynchronous temporal structure throughout the network, usually requiring specialized architectures such as Spiking Neural Networks (SNNs) (ES1).
@@ D.1 Spiking Neural Networks
Spiking Neural Networks (SNNs) are considered the third generation of neural networks, offering a biologically plausible framework for processing temporal data in a highly efficient and event-driven manner (maass1997networks). Unlike conventional artificial neural networks (ANNs) that rely on continuous-valued activations, SNNs communicate through discrete spikes, enabling sparse and asynchronous computations that closely mimic the behavior of real biological neurons. Each neuron in an SNN integrates incoming spikes over time through a membrane potential and generates an output spike only when this potential exceeds a certain threshold (gerstner2002spiking).
This temporal coding scheme allows SNNs to naturally model time-dependent dynamics, making them especially suitable for neuromorphic sensors such as event cameras, which output asynchronous streams of spatiotemporal events. Additionally, SNNs have demonstrated promising energy efficiency when deployed on neuromorphic hardware like Intel’s Loihi (davies2018loihi) or the SpiNNaker platform (furber2014spinnaker), due to their sparse activity and low-power operation. However, training SNNs remains challenging because the non-differentiable nature of spike generation complicates the use of standard backpropagation. To address this, surrogate gradient methods (neftci2019surrogate) and ANN-to-SNN conversion techniques (rueckauer2017conversion) have been widely studied, enabling SNNs to learn complex tasks with performance approaching that of traditional deep learning models.
@@ D.2 SRM-based Spiking Neuron Encoding (Extended Explanation)
In this section, we provide a detailed description of the SRM-based spiking neuron dynamics used in our SNN architecture for event stream super-resolution (SRM). Our model leverages temporal encoding and biological realism to effectively process asynchronous spike events.
@@ D.4 PSP-based Residual Upsampling Bypass
Compared to the main spiking pathway, which involves quantized spike-based transmission and nonlinear neuron thresholds, the PSP-based bypass maintains a smoother, continuous approximation of the input signal. This helps preserve fine-grained edge and texture information that may be lost during spike quantization, especially under sparse input conditions. Such a design aligns with similar principles found in residual learning (He2016ResNet), where shortcut connections assist in gradient propagation and retain low-level details.
@@ E.3 Downstream Application: Object Classification
As described in the main paper, in addition to image reconstruction, we also evaluated the performance of the super-resolved event data on an object classification task. We employed the Event Spike Tensor-based event classification method proposed by Gehrig et al. (Classification) for this evaluation. The experimental results are presented in Table 9. Since the baseline method has already achieved the highest classification accuracy, equivalent to that of original high-resolution events, on the N-MNIST and ASL-DVS datasets, our method maintains the same performance. However, on the CIFAR10-DVS dataset, our method shows a slight improvement.